OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 1 less than or greater to 9 or 2x greater than or equal to 24. A number line with an open circle on 10, shading to the left, and an open circle on 12, shading to the right A number line with an open circle on 10, an open circle on 12, and shading in between A number line with a closed circle on 10, a closed circle on 12, and shading in between A number line with a closed circle on 10, shading to the left, and a closed circle on 12, shading to the right

OpenStudy (anonymous):

@phi @ParthKohli @jigglypuff314

OpenStudy (anonymous):

@KingGeorge

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

Some body please help???

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@robtobey

jigglypuff314 (jigglypuff314):

"less than or greater to" ???

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

the less than or greater to is what confused me..

jigglypuff314 (jigglypuff314):

yeah... might be a typo if it is "less than or equal to" then D is correct :)

OpenStudy (anonymous):

thx, can you help me with one more?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 2x + 5 > 9 and 5x less than or greater to 30. A number line with an open circle on 2, shading to the left, and a closed circle on 6, shading to the right A number line with an open circle on 2, a closed circle on 6, and shading in between A number line with a closed circle on 2, shading to the left, and an open circle on 6, shading to the right A number line with a closed circle on 2, an open circle on 6, and shading in between

OpenStudy (anonymous):

Im good at everything else in math, this part in math is rly hard for me..

jigglypuff314 (jigglypuff314):

well think of it this way, 2x + 5 = 9 solve for x and 5x = 30 solve for x

OpenStudy (anonymous):

x < 2 and x≤6

jigglypuff314 (jigglypuff314):

I got x > 2 ?

OpenStudy (anonymous):

made a typo though, sorry

OpenStudy (anonymous):

ok so A?

jigglypuff314 (jigglypuff314):

x>2 means shading to the right, right? :)

OpenStudy (anonymous):

OHH, omg sorry I am so stupid, rofl

OpenStudy (anonymous):

im gonna go with B the others don't make sense to me...

jigglypuff314 (jigglypuff314):

yup, B is correct :)

OpenStudy (anonymous):

ok thanks, omg i have a few questions left and they are all confusing me... :( don't ind helping me with one or two more?

jigglypuff314 (jigglypuff314):

I don't mind :) ask away ^_^

OpenStudy (anonymous):

okay thanks, you are a life saver :)

OpenStudy (anonymous):

A company did a quality check on all the packs of nuts it manufactured. Each pack of nuts is targeted to weigh 18.25 oz. A pack must weigh within 0.36 oz of the target weight to be accepted. What is the range of rejected masses, x, for the manufactured nuts? x < 17.89 or x > 18.61 because |x – 18.25| > 0.36 x < 17.89 or x > 18.61 because |x – 0.36| + 18.25 > 0 x < 18.25 or x > 18.61 because |x – 18.25| > 0.36 x < 18.25 or x > 18.61 because |x – 0.36| + 18.25 > 0

OpenStudy (anonymous):

I have an A in my math class that I don't deserve, rofl

jigglypuff314 (jigglypuff314):

r u sure the signs are correct for the options???

OpenStudy (anonymous):

not really, the options kind of look awful to me.. :/

jigglypuff314 (jigglypuff314):

oh right nvm "rejected masses" eh anyways within 0.36 so 18.25 - 0.36 = ? and 18.25 + 0.36 = ?

OpenStudy (anonymous):

18.25 - 0.36 = 17.89 18.25 + 0.36 = 18.61

jigglypuff314 (jigglypuff314):

yes then | x - target | > (lee way)

OpenStudy (anonymous):

so either D or A

jigglypuff314 (jigglypuff314):

yeah and look at the numbers I had you calculate

OpenStudy (anonymous):

so the correct answer is A

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

thank you so much :)

jigglypuff314 (jigglypuff314):

you're welcome :) do you have more? :P

OpenStudy (anonymous):

only one, i am pretty good at the next ones after that

OpenStudy (anonymous):

Zoya has to earn at least $300 to meet her fundraising goal. She has only 100 bracelets that she plans to sell at $5 each. Which inequality shows the number of bracelets, x, Zoya can sell to meet her goal? 60 ≤ x ≤ 100 60 ≤ x ≤ 300 100 ≤ x ≤ 300 20 ≤ x ≤ 100

jigglypuff314 (jigglypuff314):

figure out the least number of bracelets she can sell 300 divided by 5 = ?

OpenStudy (anonymous):

60

jigglypuff314 (jigglypuff314):

so that will be the least amount of bracelets she can sell then now how much can she sell at max?

OpenStudy (anonymous):

100

jigglypuff314 (jigglypuff314):

there :) you have your two numbers now ^_^

OpenStudy (anonymous):

so it's A?

jigglypuff314 (jigglypuff314):

yep

OpenStudy (anonymous):

THANK YOU SO MUCH OMG i wish i could give you ten medals, lol

jigglypuff314 (jigglypuff314):

xD you're welcome :P

OpenStudy (anonymous):

:)